OpenStudy (anonymous):

xy=240, find x-y when x+y=32 and x>y can some one please provide the solution

OpenStudy (shiraz14):

Hint: Use the square of the functions.

OpenStudy (anonymous):

\[(x+y)^2=x^2+2xy+y^2=32^2\] and \((x-y)^2=x^2-2xy+y^2\)

OpenStudy (mathmale):

Focus on solving the system of equations xy=240 and x+y=32 first. Once you have x and y, separately, and make sure that x>y, you can then write x-y.